OpenStudy (juana02):

Buying a House. Sadaf Din is purchasing a house selling for $215,000. To obtain her mortgage, Sadaf must make a 20% down payment. The current mortgage rate is 4.5%. Determine the amount of the required down payment.

OpenStudy (inowalst):

@dan815

OpenStudy (juana02):

@iGreen. Can u guide me with this question?

OpenStudy (anonymous):

Find 20% of 215,000 20% = 0.2 215000 * 0.2 = ?

OpenStudy (juana02):

Ok and how do I determine the monthly principal and interest payment for a 25-year mortgage with a 20% down payment?

OpenStudy (juana02):

I got for 215000*0.2=43,000

OpenStudy (anonymous):

Yes, so that is the down payment

OpenStudy (juana02):

Ok how do I determine the principal and interest payment for the 25 year mortgage of the 43,000?

OpenStudy (juana02):

The rate is 4.5

OpenStudy (juana02):

@iGreen.

OpenStudy (anwaarullah):

p=(1+r)^25

OpenStudy (anwaarullah):

r ins interest